# Spec-Gaussian: Anisotropic View-Dependent Appearance for 3D Gaussian Splatting

### v1

@@ 3.3 Anchor-Based Gaussian Splatting
3.3.2 Adaptive Control of Anchor Gaussians.
To enable 3D-GS to represent scene details while removing redundant entities, we adaptively adjust the number of anchor Gaussians based on the gradient and opacity of the neural Gaussians. Following [21, 30], we compute the averaged gradients of the $k$ spawned neural Gaussians every 100 training iterations for each anchor Gaussian, denoted as $\nabla_{v}$. Anchor Gaussians with $\nabla_{v}>\tau_{g}$ will be densified. In practice, we follow [30] to quantize the space into multi-resolution voxels to allow new anchor Gaussians to be added at different granularities:
$\epsilon^{(l)}=\epsilon\cdot\beta/4^{l},\quad\tau_{g}^{(l)}=\tau_{g}\cdot 2^{l},$ (13)
where $l$ denotes the level of new anchor Gaussians, $\epsilon^{(l)}$ is the voxel size at the $l$-th level for newly grown anchor Gaussians, and $\beta$ represents a growth factor. Different from [30], to reduce overfitting caused by excessive densification of anchors, we introduced a hierarchical selection. Only anchor Gaussians with $\nabla_{v}>\text{Quantile}(\nabla_{v},2^{-(l+1)})$ will be densified at the corresponding voxel center at the $l$-th level.
To eliminate trivial anchors, we accumulate the opacity values of their associated neural Gaussians for every 100 training iteration, denoted as $\bar{\sigma}$. If an anchor Gaussian fails to produce neural Gaussians with a satisfactory level of opacity, with $\bar{\sigma}<\tau_{o}$, we remove it.
Dataset Mip-NeRF360 Tanks&Temples Deep Blending
Method — Metrics PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS Mem
Mip-NeRF360 27.69 0.792 0.237 0.06 8.6MB 22.22 0.759 0.257 0.14 8.6MB 29.40 0.901 0.245 0.09 8.6MB
iNGP 25.59 0.699 0.331 9.43 48MB 21.72 0.723 0.330 14.4 48MB 23.62 0.797 0.423 2.79 48MB
Plenoxels 23.08 0.626 0.463 6.79 2.1GB 21.08 0.719 0.379 13.0 2.3GB 23.06 0.795 0.510 11.2 2.7GB
3D-GS 27.47 0.812 0.222 115 748MB 23.71 0.844 0.178 169 432MB 29.65 0.899 0.247 130 662MB
Scaffold-GS 27.66 0.807 0.236 96 203MB 23.96 0.853 0.177 143 89.5MB 30.21 0.906 0.254 179 63.5MB
Ours-w/o anchor 27.81 0.810 0.223 25 1.02GB 23.94 0.846 0.181 37 563MB 29.71 0.901 0.250 32 793MB
Ours 28.01 0.812 0.222 70 245MB 24.58 0.855 0.174 111 96.5MB 30.45 0.906 0.252 132 68MB
@@
Table 1: Quantitative evaluation of our method compared to previous work on real-world datasets. We report PSNR, SSIM, LPIPS(VGG) and color each cell as best, second best and third best. Our method has overall achieved the best rendering quality, while also striking a good balance between FPS and the storage memory of 3D Gaussians.
3.4 Coarse-to-fine Training
We observed that in many real-world scenarios, 3D-GS tends to overfit the training data, leading to the emergence of numerous floaters when rendering images from novel viewpoints. A common challenge in real-world datasets is inaccuracies in camera pose estimation, particularly evident in large scenes. Scaffold-GS [30], by anchoring 3D-GS, imposes a sparse voxel constraint on the geometry, creating a hierarchical 3D-GS representation. While this hierarchical approach improves the ability of 3D-GS to model complex geometries, it does not address the overfitting issue and, in many cases, exacerbates the presence of floaters in scene backgrounds.
To mitigate the occurrence of floaters in real-world scenes, we propose a coarse-to-fine training mechanism. We believe that the tendency of 3D-GS to overfit stems from an excessive focus on each 3D Gaussian’s contribution to a specific pixel and its immediate neighbors, rather than considering broader global information. Therefore, we decide to train 3D-GS progressively from low to high resolution:
$r(i)=\min(\lfloor r_{s}+(r_{e}-r_{s})\cdot i/{\tau}\rceil,r_{e}),$ (14)
where $r(i)$ is the image resolution at the $i$-th training iteration, $r_{s}$ is the starting image resolution, $r_{e}$ is the ending image resolution (the full resolution we aim to render), and $\tau$ is the threshold iteration, empirically set to 20k.
This training approach enables 3D-GS to learn global information from the images in the early stages of training, thereby reducing overfitting to local areas of the training images and eliminating a significant number of floaters in novel view rendering. Additionally, due to the lower resolution training in the initial phase, this mechanism reduces training time by approximately 20%.
@@ 3.5 Losses
In addition to the color loss in 3D-GS [21], we also incorporate a regularization loss to encourage the neural Gaussians to remain small and minimally overlapping. Consequently, the total loss function for all learnable parameters and MLPs is formulated as:
$\mathcal{L}=(1-\lambda_{\text{D-SSIM}})\mathcal{L}_{1}$ $+\lambda_{\text{D-SSIM}}\mathcal{L}_{\text{D-SSIM}}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}},$ (15)
$\mathcal{L}_{\text{reg }}$ $=\frac{1}{N}\sum_{i=1}^{N_{\mathrm{n}}}\operatorname{Prod}\left(s_{i}\right),$
where $N_{n}$ is the number of neural Gaussians and $\operatorname{Prod}(\cdot)$ calculates the product of the scale $s_{i}$ of each neural Gaussian. The $\lambda_{\text{D-SSIM}}=0.2$ and $\lambda_{\text{reg}}=0.01$ are consistently used in our experiments.
Dataset NeRF Synthetic
Method — Metrics PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS Mem
iNGP-Base 33.18 0.963 0.045 $\sim$10 $\sim$13MB
Mip-NeRF 33.09 0.961 0.043 ¡1 $\sim$10MB
Tri-MipRF 33.65 0.963 0.042 $\sim$5 $\sim$60MB
3D-GS 33.32 0.970 0.031 415 69MB
GS-Shader 33.38 0.968 0.029 97 29MB
Scaffold-GS 33.68 0.967 0.034 240 19MB
Ours-w anchor 33.96 0.969 0.032 162 20MB
Ours 34.12 0.971 0.028 105 79MB
Table 2: Quantitative results on NeRF synthetic dataset. Our method achieves a rendering quality that surpasses NeRF-based methods, without excessively reducing FPS.
@@ 4 Experiments
In this section, we present both quantitative and qualitative results of our method. To evaluate its effectiveness, we compared it to several state-of-the-art methods across various datasets. We color each cell as best, second best and third best. Our method demonstrates superior performance in modeling complex specular and anisotropic features, as evidenced by comparisons on the NeRF, NSVF, and our ”Anisotropic Synthetic” datasets. Additionally, we showcase its versatility by comparing its performance across all scenarios in 3D-GS, further proving the robustness of our approach.
Dataset NSVF Synthetic
Method — Metrics PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS Mem
TensoRF 36.52 0.982 0.026 $\sim$1.5 $\sim$65MB
Tri-MipRF 34.58 0.973 0.030 $\sim$5 $\sim$60MB
NeuRBF 37.80 0.986 0.019 $\sim$1 $\sim$580MB
3D-GS 37.07 0.987 0.015 403 66MB
GS-Shader 33.85 0.981 0.020 68 33MB
Scaffold-GS 36.43 0.984 0.017 218 17MB
Ours-w anchor 37.71 0.987 0.015 142 18MB
Ours 38.35 0.988 0.013 91 99MB
Table 3: Quantitative results on NSVF synthetic dataset. Our method achieved significantly higher rendering quality than 3D-GS, and it also surpassed NeRF-based methods.
@@ 4.1 Implementation Details
We implemented our framework using PyTorch [37] and modified the differentiable Gaussian rasterization to include depth visualization. For the ASG appearance field, the feature decoupling MLP $\Psi$ consists of 3 layers, each with 64 hidden units, and the positional encoding for the view direction is of order 2. In terms of anchor-based Gaussian splatting, we established three levels for anchor Gaussian densification, setting the densification threshold $\tau_{g}$ to 0.0002, the pruning threshold $\tau_{o}$ to 0.005, and the number of neural Gaussians of each anchor $k$ to 10. In bounded scenes, our voxel size $\epsilon$ is set to 0.001 with a growth factor $\beta$ of 4. For the Mip-360 scenes, the voxel size remains 0.001, but the growth factor $\beta$ is increased to 16. Regarding coarse-to-fine training, we start with a resolution $r_{s}$ that is 8x downsampled. To further accelerate rendering speed, we prefilter the visible anchor Gaussians and allow only those neural Gaussians with opacity $\sigma_{n}>0$ to pass through the ASG appearance field and Gaussian rasterization pipelines. All experiments were conducted on a Tesla V100, and FPS measurements were performed on an NVIDIA RTX 3090 with 24GB of memory.
Dataset Anisotropic Synthetic
Method — Metrics PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS Mem
3D-GS 33.82 0.966 0.062 345 47MB
Scaffold-GS 35.34 0.972 0.052 234 27MB
Ours-w anchor 36.76 0.976 0.046 180 28MB
Ours 37.42 0.977 0.047 119 59MB
Table 4: Quantitative results on our ”Anisotropic Synthetic” dataset.
Refer to caption Refer to caption Refer to caption Refer to caption
Scaffold-GS ASG color Ours GT
Figure 5: Ablation on ASG feature decoupling MLP. We show that directly using ASG to model color leads to the failure in modeling anisotropy and specular highlights. By decoupling the ASG features through MLP, we can realistically model complex optical phenomena.
@@ 4.2 Results and Comparisons
We used the NeRF, NSVF, and our ”Anisotropic Synthetic” datasets as the experimental datasets for synthetic scenes. Our comparisons were made with the most relevant state-of-the-art methods, including 3D-GS [21], Scaffold-GS [30], GaussianShader [20], and several NeRF-based methods such as NSVF [28], TensoRF [6], NeuRBF [9], and Tri-MipRF [17]. To ensure a fair comparison, we used the rendering metrics of the NeRF and NSVF datasets as reported in the baseline papers. For scenes not reported in the baseline papers, we trained the baselines from scratch using the released codes and their default configurations.
As shown in Fig. 4, Figs. 7-8. and Tabs. 2-4, our method achieved the highest performance in terms of PSNR, SSIM, and LPIPS. It also significantly improved upon the issues that 3D-GS faced in modeling high-frequency specular highlights and complex anisotropy. See per-scene results in the supplementary materials.
@@
To verify the versatility of our method in real-world scenarios, we used the same real-world dataset as in 3D-GS [21]. As shown in Tab. 1, our method achieves rendering results comparable to state-of-the-art methods on the Deep Blending dataset and surpasses them on Mip-NeRF 360 and Tanks&Temples. Furthermore, our method effectively balances FPS, storage overhead, and rendering quality. It enhances rendering quality without excessively increasing storage requirements or significantly reducing FPS. As illustrated in Fig. 6, our method has also significantly improved the visual effect. It removes a large number of floaters in outdoor scenes and successfully models the high-frequency specular highlights in indoor scenes. This demonstrates that our approach is not only adept at modeling complex specular scenes but also effectively improves rendering quality in general scenarios.
@@ 4.3 Ablation Study
4.3.1 ASG feature decoupling MLP
We conducted an ablation study to evaluate the effectiveness of using ASG to output features, which are then decoupled through an MLP $\Psi$ to derive the final specular color. As demonstrated in Fig. 5, directly using ASG to output color results in the inability to model specular and anisotropic components. In contrast to directly using an MLP for color modeling, as in Scaffold-GS [30], ASG can encode higher-frequency anisotropic features. This capability aids the MLP in learning complex optical phenomena, leading to more accurate and detailed rendering results.
4.3.2 Coarse-to-fine training
We conducted an ablation study to assess the impact of coarse-to-fine (c2f) training. As illustrated in Fig. 10, both 3D-GS and Scaffold-GS exhibit a large number of floaters in the novel view synthesis. Coarse-to-fine training effectively reduces the number of floaters, alleviating the overfitting issue commonly encountered by 3D-GS in real-world scenarios.
@@ 5 Conclusion
In this work, we introduce Spec-Gaussian, a novel approach to 3D Gaussian splitting that features an anisotropic view-dependent appearance. Leveraging the powerful capabilities of ASG, our method effectively overcomes the challenges encountered by vanilla 3D-GS in rendering scenes with specular highlights and anisotropy. Additionally, we innovatively implement a coarse-to-fine training mechanism to eliminate floaters in real-world scenes. Both quantitative and qualitative experiments demonstrate that our method not only equips 3D-GS with the ability to model specular highlights and anisotropy but also enhances the overall rendering quality of 3D-GS in general scenes, without significantly compromising FPS and storage overhead.
Limitations
Although our method enables 3D-GS to model complex specular and anisotropic features, it still faces challenges in handling reflections. Specular and anisotropic effects are primarily influenced by material properties, whereas reflections are closely related to the environment and geometry. Due to the lack of explicit geometry in 3D-GS, we cannot differentiate between reflections and material textures using constraints like normals, as employed in Ref-NeRF [45] and NeRO [29]. In our experiments, we also observed that when ground truth geometric information is provided, 3D-GS becomes more consistent with expectations under strict constraints, but this comes at the cost of a certain decline in rendering quality. We plan to explore solutions for modeling reflections with 3D-GS in future work.
Appendix A More Results

### v2

@@ 3.4 Adaption for Anchor-Based Gaussian Splatting
Dataset Mip-NeRF 360 Mip-NeRF 360 Outdoor Mip-NeRF 360 Indoor
Method | Metrics PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS Mem PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Plenoxels 23.08 0.626 0.463 6.79 2.1GB 21.68 0.513 0.491 24.83 0.766 0.426
iNGP 25.59 0.699 0.331 9.43 48MB 22.75 0.567 0.403 29.14 0.863 0.242
Mip-NeRF360 27.69 0.792 0.237 0.06 8.6MB 24.47 0.691 0.283 31.72 0.917 0.180
3D-GS 27.79 0.826 0.202 115 748MB 25.02 0.742 0.232 31.25 0.931 0.164
Scaffold-GS 27.98 0.824 0.207 96 203MB 25.07 0.736 0.243 31.61 0.933 0.162
Ours-w/ anchor 28.14 0.824 0.196 70 260MB 24.98 0.735 0.223 32.09 0.935 0.161
Ours-light 28.07 0.834 0.183 44 684MB 25.09 0.752 0.203 31.80 0.936 0.158
Ours 28.18 0.835 0.176 33 847MB 25.11 0.754 0.195 32.01 0.937 0.153
@@
Table 1: Quantitative comparison of on real-world datasets. We report PSNR, SSIM, LPIPS (VGG) and color each cell as best, second best and third best. Our method has achieved the best rendering quality, while striking a good balance between FPS and the storage memory.
[…]
[7.8cm] Dataset NeRF Synthetic Method | Metrics PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS Mem iNGP-Base 33.18 0.963 0.045 $\sim$10 13MB Mip-NeRF 33.09 0.961 0.043 <1 10MB Tri-MipRF 33.65 0.963 0.042 $\sim$5 60MB 3D-GS 33.32 0.969 0.031 315 69MB GS-Shader 33.38 0.968 0.030 97 29MB Scaffold-GS 33.68 0.967 0.034 240 19MB Ours-w/ anchor 33.96 0.969 0.032 172 19MB Ours-light 34.08 0.970 0.029 148 58MB Ours 34.19 0.971 0.028 121 72MB
Figure 6: Results on NeRF synthetic dataset.
[…]
[7.8cm] Dataset NSVF Synthetic Method | Metrics PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS Mem TensoRF 36.52 0.982 0.026 1.5 65MB Tri-MipRF 34.58 0.973 0.030 $\sim$5 60MB NeuRBF 37.80 0.986 0.019 $\sim$1 580MB 3D-GS 37.07 0.987 0.015 306 71MB GS-Shader 33.85 0.981 0.020 68 33MB Scaffold-GS 36.43 0.984 0.017 218 17MB Ours-w/ anchor 37.71 0.987 0.015 152 16MB Ours-light 38.28 0.988 0.013 124 70MB Ours 38.40 0.988 0.012 108 89MB
Figure 7: Results on NSVF synthetic dataset.
@@ 3.5 Losses
We optimize the learnable parameters and MLPs using the same loss function as 3D-GS [20]. The total supervision is given by:
$\mathcal{L}=(1-\lambda_{\text{D-SSIM}})\mathcal{L}_{1}$ $+\lambda_{\text{D-SSIM}}\mathcal{L}_{\text{D-SSIM}},$ (14)
where the $\lambda_{\text{D-SSIM}}=0.2$ is consistently used in our experiments.
@@ 4 Experiments
In this section, we present both quantitative and qualitative results of our method. To evaluate its effectiveness, we compared it to several state-of-the-art methods across various datasets. We color each cell as best, second best and third best. Our method includes three versions, each based on different foundational methods with distinct hyperparameter settings. The performance version (Ours) is based on 3D-GS [20] with $\tau_{g}=0.0005$; the light version (Ours-light), also based on 3D-GS, has $\tau_{g}=0.0006$; and the mini version (Ours-w/ anchor) is based on Scaffold-GS [29], with $\tau_{g}=0.0006$. Our method demonstrates superior performance in modeling complex specular and anisotropic features, as evidenced by comparisons on the NeRF, NSVF, and our "Anisotropic Synthetic" datasets. Additionally, we showcase its versatility by comparing its performance in diffuse scenarios, further proving the robustness of our approach.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Mip-NeRF 360 3D-GS Scaffold-GS Ours GT
Figure 8: Visualization on Mip-NeRF 360 indoor scenes. Our method achieves superior recovery of specular effects compared to SOTA methods.
@@ 4.1 Implementation Details
We implemented our framework using PyTorch [39] and modified the differentiable Gaussian rasterization to include depth visualization. For the ASG appearance field, the decoupling MLP $\Psi$ consists of 3 layers, each with 64 hidden units, and the positional encoding for the view direction is of order 2. Regarding coarse-to-fine training, which is applied only to real-world scenes to remove floaters, we start with a resolution $r_{s}$ that is 4x downsampled. To further accelerate rendering, we prefilter and allow only those Gaussians with opacity $\sigma_{n}>0$ to pass through the ASG appearance field and Gaussian rasterization pipelines. All experiments were conducted on an NVIDIA RTX 3090.
@@ 4.2 Results and Comparisons
We used the NeRF, NSVF, and our "Anisotropic Synthetic" datasets as the experimental datasets for synthetic scenes. Our comparisons were made with the most relevant state-of-the-art methods, including 3D-GS [20], Scaffold-GS [29], GaussianShader [19], and several NeRF-based methods such as NSVF [27], TensoRF [5], NeuRBF [8], and Tri-MipRF [16].
As shown in Fig. 5 (with PSNR and LPIPS), and Tabs. 77, our method achieved the highest performance with fewer Gaussians compared to vanilla 3D-GS. It also improved upon the issues that 3D-GS faced in modeling high-frequency specular highlights and complex anisotropy as shown in Tab. 4 with fewer Gaussians and better metrics. See more in the supplementary materials.
@@
To verify the versatility of our method in real-world scenarios, we used the Mip360 [2] dataset, which contains indoor scenes with specular highlights. As shown in Tab. 1, our method surpasses state-of-the-art methods on Mip-NeRF 360. Furthermore, our method effectively balances FPS, storage, and rendering quality. It enhances rendering quality without increasing storage or significantly reducing FPS. As illustrated in Fig. 8 and Fig. 10, our method has also significantly improved the visual effect. It removes a large number of floaters in outdoor scenes and successfully models the high-frequency specular highlights in indoor scenes. This demonstrates that our approach is not only adept at modeling complex specular scenes but also effectively improves rendering quality in general scenarios.
Figure 9: Ablation on ASG appearance field. We show that directly using ASG to model color leads to the failure in modeling anisotropy and specular highlights. By decoupling the ASG features through MLP, we can realistically model complex optical phenomena.
@@ 4.3 Ablation Study
ASG feature decoupling MLP.
We conducted an ablation study to evaluate the key components of the ASG appearance field, which include ASG features, decoupling MLP, and the separation of diffuse and specular colors. As demonstrated in Fig. 9 (with PSNR and LPIPS), directly using ASG to output color results in the inability to model specular and anisotropic components. In contrast to directly using an MLP for color modeling, as in Scaffold-GS [29], separately modeling diffuse and specular color can enhance the fitting ability for high-frequency information. ASG can encode higher-frequency anisotropic features. With the help of ASG’s ability to encode high-frequency anisotropic features, the decoupling MLP can fit complex optical phenomena, leading to more accurate rendering results. We also demonstrated that higher-order SH (6-order) and more MLP layers (4-layers) do not help 3D-GS and Scaffold-GS achieve satisfactory results, highlighting the importance of ASG.
Coarse-to-fine training.
We conducted an ablation study to assess the impact of coarse-to-fine (c2f) training. As illustrated in Fig. 10 (with LPIPS and number of Gaussian), both 3D-GS and Scaffold-GS exhibit a large number of floaters in the novel view synthesis. Coarse-to-fine training effectively reduces the number of floaters, alleviating the overfitting issue commonly encountered by 3D-GS in real-world scenarios. Applying an L1 constraint to the gradients used for 3D-GS densification further reduced the number of floaters and Gaussians. See more in the supplementary materials.
Figure 10: Ablation on coarse-to-fine training. Experimental results demonstrate that our simple yet effective training mechanism can effectively remove floaters without increasing the number of 3D Gaussians, thereby alleviating the overfitting problem prevalent in 3D-GS-based methods.
@@ 5 Conclusion
In this work, we introduce Spec-Gaussian, a novel approach to 3D Gaussian splitting that features an anisotropic view-dependent appearance. Leveraging the powerful capabilities of ASG, our method effectively overcomes the challenges encountered by vanilla 3D-GS in rendering scenes with specular highlights and anisotropy. Additionally, we innovatively implement a coarse-to-fine training mechanism to eliminate floaters in real-world scenes. Both quantitative and qualitative experiments demonstrate that our method not only equips 3D-GS with the ability to model specular highlights and anisotropy but also enhances the overall rendering quality of 3D-GS in general scenes.
Limitations.
Although our method enables 3D-GS to model complex specular and anisotropic features, it still faces challenges in handling reflections. Specular and anisotropic effects are primarily influenced by material properties, whereas reflections are closely related to the environment and geometry. Due to the lack of explicit geometry in 3D-GS, we cannot differentiate between reflections and materials using constraints like normals, as employed in Ref-NeRF [47] and NeRO [28]. We plan to explore solutions for modeling reflections with 3D-GS in future work.
6 Acknowlegements
We thank Chao Wan from Cornell University for the help during rebuttal period. This work was supported by the National Natural Science Foundation of China (Grant No. 62036010). Ziyi Yang was also supported by ByteDance MMLab.
Appendix A Appendix / supplemental material
This supplementary material provides more results that accompany the paper.
  • Section B provides more ablations.
  • Section C provides additional results, including more visualizations and quantitative results on complete datasets.
Appendix B More Ablations
In this section, we present the complete quantitative ablations on the key components of our method.
We first evaluate the role of each component of the ASG appearance field in NeRF synthetic scenes as shown in Tab. 12. The introduction of ASG improves the ability to model specular highlights and reduces the number of 3D Gaussians. The inclusion of normals did not significantly increase computational overhead, but it did enhance rendering metrics and visual quality. More importantly, we achieve better rendering quality with fewer Gaussians than vanilla 3D-GS, a characteristic that can be further explored in the future.
Next, we evaluated our full method on the Mip360 dataset in Tab. 12. It is important to note that the Mip360 dataset is divided into indoor and outdoor scenes. Indoor scenes have more specular highlights, while outdoor scenes contain a large number of floaters. The coarse-to-fine approach itself improves the quality of 3D-GS in real-world scenes, mainly by eliminating a significant amount of floaters in outdoor settings. Although the introduction of the ASG appearance field significantly increases rendering overhead, it did greatly enhance the modeling of specular highlights in indoor scenes. Under the constraints of the coarse-to-fine mechanism, our complete method combines the advantages of both, achieving the best rendering quality. To further improve rendering speed, we also implement a light version and a mini version based on Scaffold-GS. These versions offer a trade-off between rendering quality and speed and can be used as needed. The quality of the Mip360 scenes demonstrates that our method is not only capable of handling scenes with specular highlights but is also robust in real-world diffuse scenarios.
Appendix C More Comparisons